OpenStudy (anonymous):

Write the equation in polar coordinates: x^2+y^2=4

OpenStudy (anonymous):

@satellite73 here

OpenStudy (anonymous):

@saifoo.khan here

OpenStudy (anonymous):

hi :D

OpenStudy (anonymous):

circle with center (0,0) radius fixed at 2 right?

OpenStudy (anonymous):

radius is right

OpenStudy (anonymous):

so \(\theta\) can be anything but \(r=2\)

OpenStudy (anonymous):

the answer just radius, nothing else, i dont understand lol o_0

OpenStudy (anonymous):

so x^2 and y^2 = 0,0 ?

OpenStudy (anonymous):

yes you are fixing the radius at \(r=2\) whereas \(\theta\) can take on any value

OpenStudy (anonymous):

equation is \(r=2\)

OpenStudy (anonymous):

so what if it was x^2+y^2=9 then it would be (0,0) r=3 ?

OpenStudy (anonymous):

think by analogy to the equation for a horizontal line in rectangular (regular old) coordinates. say at \(y=2\) then we just write \(y=2\) without the \(x\) because \(x\)can be anything, \(y\) does not depend on \(x\), and \(y\) is a fixed number just like in your example above, \(\theta\) can take on any value, but \(r=2\) a fixed number

OpenStudy (anonymous):

yes you are right